# Chemical Abundances in the Leiptr Stellar Stream: a Disrupted Ultra-faint Dwarf Galaxy?

### v1

@@ 3.1 Differential Abundance Analysis
We thus conduct a differential abundance analysis between these two pairs of stars to examine whether a more precise analysis can distinguish between the star compositions. Adopting the same stellar parameters as our main analysis, we calculate abundance differences for matched lines between the two pairs of stars using Table 3111Note that differential abundance analyses in the literature usually use one reference star to compare to all other stars, but in this case we are simply looking at two pairs of stars with similar stellar parameters.. We estimate uncertainties by taking the standard deviation and standard error of the line abundance differences, and not propagating any stellar parameter uncertainties that we expect to largely cancel out in this differential analysis.
@@ 4 Abundance Results
We first discuss the overall metallicity dispersion in the Leiptr stream in Section 4.1. Then, we discuss the abundance trends for halo, dwarf galaxy, and globular cluster stars in each element group (Sections 4.2-4.6) and compare to our results for Leiptr. Our literature sample is compiled based on references whose analysis process is very similar to ours (i.e., high-resolution optical spectroscopy of individual stars, mostly with spectroscopic stellar parameters). Abundances for stars in the metal-poor Milky Way halo come from selected references in JINAbase (Abohalima & Frebel 2018, from Barklem et al. 2005; Cohen et al. 2013; Roederer et al. 2014; Jacobson et al. 2015) and abundances of GSE stars come from Carrillo et al. (2022). We take stellar masses of dwarf galaxies from the compilation by Simon (2019). Massive dwarf galaxy () abundances come from Lemasle et al. 2014 (Fornax), Jönsson et al. 2020 (Sagittarius, APOGEE stars), and Hill et al. 2019 (Sculptor). Intermediate-mass dwarf galaxy () abundances come from Norris et al. 2017 (Carina), Cohen 2009 (Draco), Theler et al. 2020 (Sextans), and Cohen & Huang 2010 (Ursa Minor). Ultra-faint dwarf galaxy () abundances come from Koch et al. (2008); Feltzing et al. (2009); Norris et al. (2010b, a); Simon et al. (2010); Frebel et al. (2010, 2014); Gilmore et al. (2013); Koch et al. (2013); Roederer & Kirby (2014); Ishigaki et al. (2014); François et al. (2016); Ji et al. (2016c, b, 2019, 2020b); Venn et al. (2017); Kirby et al. (2017); Hansen et al. (2017, 2020b, 2024); Nagasawa et al. (2018); Spite et al. (2018); Marshall et al. (2019); Chiti et al. (2018, 2023); Waller et al. (2023); and Webber et al. (2023). Chemical abundances for Milky Way globular clusters come from Carretta et al. (2009), Cohen & Kirby (2012), Kirby et al. (2023), Roederer et al. (2011), and Worley et al. (2013). Stars from dwarf galaxy and globular cluster streams come from Ji et al. (2020a).
@@
The four Leiptr member stars fall into two categories. The three lowest-metallicity stars (Leiptr-186, Leiptr-208 and Leiptr-342) have similar element abundances: low , odd-z, and neutron-capture ratios. The most metal-rich star (Leiptr-321) has similar abundances to the low-metallicity stars, but higher and odd-z ratios. We further investigate our pairs from Section 3.1 by marking Leiptr-208 and Leiptr-321 as circles and Leiptr-186 and Leiptr-342 as squares in Figures 7-11. Halo stars are marked as plus signs in gray if from the Milky Way or dark-blue if from GSE; dwarf galaxy stars are marked as triangles in yellow if massive, pink if intermediate-mass, or purple if ultra-faint; globular clusters are marked as upside-down triangles in dark-red; and stellar streams are marked as stars in pink if from a dwarf galaxy or dark-red if from a globular cluster.
@@ 4.1 Metallicity Mean and Dispersion
The four Leiptr stars fall between [Fe/H] = and . We follow the method used in Usman et al. (2024) to determine the mean metallicity and intrinsic dispersion, assuming a single Gaussian with heteroskedastic uncertainties, sampled using the nested sampling algorithm dynesty (Speagle, 2020; Higson et al., 2019; Koposov et al., 2022) to calculate the Bayesian posteriors of the means and dispersions. The reported uncertainties are determined with 1 percentiles and 90% upper limits.
Using our full abundance uncertainties, which are dominated by propagating both statistical and systematic uncertainties in stellar parameters, the mean [Fe I/H] metallicity of the four Leiptr stars is , with a dispersion 90% upper limit less than 0.38. If instead we use [Fe II/H], the per-star uncertainties are reduced to dex for each star, as the effect of our surface gravity uncertainties on Fe II are much less than the effect of our effective temperature uncertainties on Fe I. We then resolve a [Fe II/H] spread to be 0.29, with 90% confidence that the spread is larger than 0.1 dex. We note the mean metallicity is 0.6 dex lower than the photometric estimate by Ibata et al. (2019), indicating that broad-band photometric metallicities are not sufficiently accurate to be taken at face value.
A resolved metallicity dispersion, indicating internal enrichment by supernovae, is crucial for classifying Leiptr as a dwarf galaxy (Willman & Strader, 2012). We thus performed a more sensitive test of internal metallicity variations by looking at our pairwise differential abundances (Section 3.1,Figure 6). These clearly show that our full uncertainties for [Fe I/H] are too conservative for intrinsic dispersion measurements. The four likely Leiptr member stars come in two pairs of stars with similar stellar parameters and clearly distinct chemical abundances. Thus, the metallicity differences and [Fe II/H] metallicity dispersion strongly suggest that Leiptr is a dwarf galaxy, rather than a globular cluster.
@@ 4.2 C and N
In principle, carbon and nitrogen abundances can be used to distinguish between globular clusters and dwarf galaxies. This is because the majority of stars in massive globular clusters display unusual carbon depletions and nitrogen enhancements, a phenomenon whose origin is not understood but is empirically well-documented (e.g., Bastian & Lardo, 2018; Gratton et al., 2019). We use the 4310 and 4323 Å CH absorption bands to determine the carbon abundance and 3877 Å CN band to find upper limits on nitrogen for Leiptr-186, Leiptr-208, and Leiptr-321. We find that Leiptr-342 has high nitrogen and low carbon. Thus, this star could potentially indicate that Leiptr is a disrupted globular cluster.
However, red giant branch stars also change composition as they ascend the red giant branch due to extra mixing that also depletes carbon and enhances nitrogen (e.g., Shetrone et al., 2019), so we must also examine the relationship between chemical abundance and effective temperature. We plot carbon and nitrogen abundance against in the right two panels of Figure 7. We expect to see higher nitrogen and lower carbon as decreases due to extra mixing (e.g., Figure 6 of Spite et al. 2005). Enriched globular cluster stars would not follow this temperature dependence and instead show a distinct group of stars with high nitrogen and low carbon, irrespective of measured temperature. Compared to Cohen et al. (2005), which specifically targeted stars in M15 that are on the lower red giant branch to avoid extra mixing, the Leiptr stars are cooler and within the regime of possible extra mixing effects. Since we do observe that the coolest Leiptr stars have higher nitrogen and lower carbon, the trend with suggests that extra mixing is most likely the reason for Leiptr’s chemical abundance pattern. Furthermore, the nitrogen increase is not as extreme as it is in other globular clusters like M15 (Cohen et al., 2005). However, with only one nitrogen determination in the Leiptr member stars, more stars would be needed for this to be a strong constraint on whether Leiptr’s progenitor is a globular cluster or dwarf galaxy.
@@
*approximate NLTE corrections are calculated as an average for aluminum and independently for sodium and potassium (see text for details)
The [/Fe] ratio is an excellent indicator of a stream’s progenitor due to galactic chemical evolution. elements are primarily produced by core-collapse supernovae, while iron is produced by both core-collapse and Type Ia supernovae. The [/Fe] ratio thus starts high early then declines to lower values later as a galaxy forms. Efficiently star-forming galaxies like the Milky Way maintain a high [/Fe] until relatively high metallicities, while inefficiently star-forming dwarf galaxies have an downturn at lower metallicities (e.g., Tinsley, 1979; Matteucci & Brocato, 1990; McWilliam, 1997; Tolstoy et al., 2009; Sheffield et al., 2012). This downturn, often called an “knee”, tends to occur at lower metallicities in lower-mass galaxies (e.g., Tolstoy et al., 2009; Kirby et al., 2011, 2020; Frebel & Bromm, 2012; Helmi et al., 2018; Hill et al., 2019; Theler et al., 2020). Globular clusters can be understood in this context as inheriting the chemical patterns of their host galaxies at their birth metallicity, as they are just the most massive star clusters to form in a galaxy (e.g., Forbes et al., 2018; Choksi et al., 2018; Reina-Campos et al., 2019). Thus, globular clusters born in the Milky Way will tend to be high-, while those born in dwarf galaxies will be low- (e.g., Mucciarelli et al., 2017; Usman et al., 2024).
Figure 8 shows that Leiptr’s stars are overall low-. While Milky Way halo stars and stars in Milky Way globular clusters M92 and NGC 2419 have [/Fe] (NGC 2419 has some low-magnesium stars, but this is from magnesium depletion due to multiple populations), the three lowest-metallicity Leiptr stars’ magnesium, calcium, and titanium abundances are within of [/Fe] = 0. hus, the Milky Way halo and its globular clusters’ high- abundances are overall inconsistent with Leiptr. However, several dwarf galaxies have low [/Fe] ratios similar to Leiptr. At , lower-mass classical dwarf galaxies () like Draco, Ursa Minor, and Sextans, as well as the dwarf galaxy stream Elqui, have low [/Fe] abundances similar to Leiptr. The lowest-mass ultra-faint dwarf galaxies like Carina II, Reticulum II, and Horologium I () also mostly have low [/Fe] abundances at . In contrast, higher-mass dwarf galaxies like Sculptor, Fornax, and GSE () still have relatively high [/Fe] ratios at , not seeing a significant decline until higher metallicity, . Thus, Leiptr’s element abundances are most similar to lower-mass dwarf galaxies.
@@ 4.4 Odd-Z elements (Na, Al, K, Sc)
*approximate NLTE corrections are calculated independently for manganese (see text for details)
Figure 9 shows the abundances for the odd-Z elements. In light-yellow points, we show the effect of applying corrections to sodium, aluminum, and potassium for NLTE effects on the lines where the radiation field is partially decoupled from local conditions. Using the calculations for neutral sodium from Lind et al. (2011), we compute the average of the corrections from each of the used lines for each star. All five stars include the 5895 Å sodium line and some of the five also use the 5682, 5688, and 5889 Å lines. The typical size of the correction is dex. Aluminum is determined using the 3961 Å resonance line, and we adopt a constant correction for all stars of adding 0.6 to [Al/Fe], following Roederer & Gnedin (2019). For potassium, we use the two resonance lines at 7664 Å and 7669 Å. The two lines always agree excellently in abundance, and we note the 7664 Å line is not affected by tellurics given Leiptr’s radial velocity (Figure 5). We identify stars in Table A.1 of Reggiani et al. (2019) that have the most comparable stellar parameters as the Leiptr stars and use the corresponding potassium NLTE corrections. We use the corrections from stars HD 186478, CS 29513-014, BD +19 1185A, HD 45282, and HD 126238 with a range from to dex. The LTE and NLTE abundances are both reported in Tables 3-4.2.
In Figure 9, compared to halo stars (whose abundances are mostly computed in LTE and should be compared to the light-blue points), Leiptr lies on the lower end of the abundance scatter for sodium, potassium, and scandium. Abundances of aluminum in the Milky Way globular clusters, M92 and NGC 2419, are also substantially higher than in all four Leiptr stars. The globular clusters should be compared to the NLTE aluminum abundances (light-yellow points): the M92 stars use the same aluminum 3961 Å line and were corrected using Nordlander & Lind (2017) while the NGC 2419 stars use the subordinate doublet at Å that are minimally affected by NLTE. These aluminum abundances further disprove that Leiptr may be an accreted globular cluster with multiple populations, which should result in aluminum enhancements as shown by the globular clusters in Figure 9 (see Bastian & Lardo, 2018; Gratton et al., 2019, for details).
@@ 4.5 Iron-peak elements (Cr, Mn, Co, Ni, Zn)
Figure 10 shows some of the iron-peak abundances in Leiptr. Manganese deserves special discussion due to NLTE effects. We reject the resonance lines at 4030 Å and only use weaker lines around 4800 Å. This was motivated due to expecting relatively small NLTE corrections (Bergemann & Gehren, 2008), but more recent calculations show that the corrections for the weaker lines can also be dex (Bergemann et al., 2019), so we apply an individual NLTE correction from Bergemann et al. (2019) to each star from the average of the corrections for each line, which appear as light-yellow points in Figure 10. We note that the comparison stars in Figure 10 do not have NLTE corrections applied, so these should be compared to the LTE abundances.

### v2

@@ 3.1 Differential Abundance Analysis
We thus conduct a differential abundance analysis between these two pairs of stars to examine whether a more precise analysis can distinguish between the star compositions. Adopting the same stellar parameters as our main analysis, we calculate abundance differences for matched lines between the two pairs of stars using Table 3111Note that differential abundance analyses in the literature usually use one reference star to compare to all other stars, but in this case we are simply looking at two pairs of stars with similar stellar parameters.. Unlike the initial abundance analysis, we do not apply any weighting for the differential analysis. We estimate uncertainties by taking the standard deviation and standard error of the line abundance differences, and not propagating any stellar parameter uncertainties that we expect to largely cancel out in this differential analysis.
@@ 4 Abundance Results
We first discuss the overall metallicity dispersion in the Leiptr stream in Section 4.1. Then, we discuss the abundance trends for halo, dwarf galaxy, and globular cluster stars in each element group (Sections 4.2-4.6) and compare to our results for Leiptr. Total uncertainty is shown in black for accuracy and abundance uncertainty without additional systematic errors is shown in the same color as the stars for statistical precision as described in Section 3. Our literature sample is compiled based on references whose analysis process is very similar to ours (i.e., high-resolution optical spectroscopy of individual stars, mostly with spectroscopic stellar parameters). Abundances for stars in the metal-poor Milky Way halo come from selected references in JINAbase (Abohalima & Frebel 2018, from Barklem et al. 2005; Cohen et al. 2013; Roederer et al. 2014; Jacobson et al. 2015) and abundances of GSE stars come from Carrillo et al. (2022). We take stellar masses of dwarf galaxies from the compilation by Simon (2019). Massive dwarf galaxy () abundances come from Lemasle et al. 2014 (Fornax), Jönsson et al. 2020 (Sagittarius, APOGEE stars), and Hill et al. 2019 (Sculptor). Intermediate-mass dwarf galaxy () abundances come from Norris et al. 2017 (Carina), Cohen 2009 (Draco), Theler et al. 2020 (Sextans), and Cohen & Huang 2010 (Ursa Minor). Ultra-faint dwarf galaxy () abundances come from Koch et al. (2008); Feltzing et al. (2009); Norris et al. (2010a, b); Simon et al. (2010); Frebel et al. (2010, 2014); Gilmore et al. (2013); Koch et al. (2013); Roederer & Kirby (2014); Ishigaki et al. (2014); François et al. (2016); Ji et al. (2016b, c, 2019, 2020b); Venn et al. (2017); Kirby et al. (2017); Hansen et al. (2017, 2020a, 2024); Nagasawa et al. (2018); Spite et al. (2018); Marshall et al. (2019); Chiti et al. (2018, 2023); Waller et al. (2023); and Webber et al. (2023). Chemical abundances for Milky Way globular clusters come from Carretta et al. (2009), Cohen & Kirby (2012), Kirby et al. (2023), Roederer et al. (2011), and Worley et al. (2013). Stars from dwarf galaxy and globular cluster streams come from Ji et al. (2020a).
@@
The four Leiptr member stars fall into two categories. The three lowest-metallicity stars (Leiptr-186, Leiptr-208 and Leiptr-342) have similar element abundances: low , odd-z, and neutron-capture ratios. The most metal-rich star (Leiptr-321) has similar abundances to the low-metallicity stars, but higher and odd-z ratios. We further investigate our pairs from Section 3.1 by marking Leiptr-208 and Leiptr-321 as circles and Leiptr-186 and Leiptr-342 as squares in Figures 7-11. Halo stars are marked as plus signs in gray if from the Milky Way or dark-blue if from GSE; dwarf galaxy stars are marked as triangles in yellow if massive ( > ), pink if intermediate-mass ( < < ), or purple if ultra-faint ( < ); globular clusters are marked as upside-down triangles in dark-red; and stellar streams are marked as stars in pink if from a dwarf galaxy or dark-red if from a globular cluster.
@@ 4.1 Metallicity Mean and Dispersion
[…]
The four Leiptr stars fall between [Fe/H] = and . We follow the method used in Usman et al. (2024) to determine the mean metallicity and intrinsic dispersion, assuming a single Gaussian with heteroskedastic uncertainties, sampled using the nested sampling algorithm dynesty (Speagle, 2020; Higson et al., 2019; Koposov et al., 2022) to calculate the Bayesian posteriors of the means and dispersions. The reported uncertainties are determined with 1 percentiles and 90% upper limits. We acknowledge that [Fe/H] may be underestimated by about 0.2 dex by ignoring NLTE effects on the measured iron lines (e.g., Bergemann et al., 2012c, a; Ezzeddine et al., 2017), but each star will shift a similar amount, so as a whole Leiptr will maintain the same metallicity dispersion.
Using our full abundance uncertainties, which are dominated by propagating both statistical and systematic uncertainties in stellar parameters, the mean [Fe I/H] metallicity of the four Leiptr stars is , with a dispersion 90% upper limit less than 0.38. If instead we use [Fe II/H], the per-star uncertainties are reduced to dex for each star, as the effect of our surface gravity uncertainties on Fe II are much less than the effect of our effective temperature uncertainties on Fe I. We then resolve a [Fe II/H] spread to be 0.29, with 90% confidence that the spread is larger than 0.1 dex. We note the mean metallicity is 0.6 dex lower than the photometric estimate by Ibata et al. (2019).
A resolved metallicity dispersion, indicating internal enrichment by supernovae, is crucial for classifying Leiptr as a dwarf galaxy (Willman & Strader, 2012). We thus performed a more sensitive test of internal metallicity variations by looking at our pairwise differential abundances (Section 3.1, Figure 6). These clearly show that our full uncertainties for [Fe I/H] are too conservative for intrinsic dispersion measurements. The four likely Leiptr member stars come in two pairs of stars with similar stellar parameters and clearly distinct chemical abundances. Thus, the metallicity differences and [Fe II/H] metallicity dispersion strongly suggest that Leiptr was a dwarf galaxy, rather than a globular cluster.
@@ 4.2 C and N
In principle, carbon and nitrogen abundances can be used to distinguish between globular clusters and dwarf galaxies. This is because the majority of stars in massive globular clusters display unusual carbon depletions and nitrogen enhancements, a phenomenon whose origin is not understood but is empirically well-documented (e.g., Bastian & Lardo, 2018; Gratton et al., 2019). We use the 4310 and 4323 Å CH absorption bands to determine the carbon abundance and 3877 Å CN band to find upper limits on nitrogen for Leiptr-186, Leiptr-208, and Leiptr-321. The NLTE correction is negligible for carbon (Amarsi et al., 2019a, b) and nitrogen (Amarsi et al., 2020; Mashonkina & Ryabchikova, 2024). We find that Leiptr-342 has high nitrogen and low carbon. Thus, this star could potentially indicate that Leiptr is a disrupted globular cluster.
However, red giant branch stars also change composition as they ascend the red giant branch due to extra mixing that also depletes carbon and enhances nitrogen (e.g., Shetrone et al., 2019), so we must also examine the relationship between chemical abundance and effective temperature. We plot carbon and nitrogen abundance against in the right two panels of Figure 7. We expect to see higher nitrogen and lower carbon as decreases due to extra mixing (e.g., Figure 6 of Spite et al. 2005). Enriched globular cluster stars would not follow this temperature dependence and instead show a distinct group of stars with high nitrogen and low carbon, irrespective of measured temperature. Compared to Cohen et al. (2005), which specifically targeted stars in M15 that are on the lower red giant branch to avoid extra mixing, the Leiptr stars are cooler and within the regime of possible extra mixing effects. Since we do observe that the coolest Leiptr stars have higher nitrogen and lower carbon, the trend with suggests that extra mixing is most likely the reason for Leiptr’s chemical abundance pattern. Furthermore, the nitrogen increase is not as extreme as it is in globular clusters like M15 (Cohen et al., 2005). However, with only one nitrogen determination in the Leiptr member stars, more stars would be needed for this to be a strong constraint on whether Leiptr’s progenitor is a globular cluster or dwarf galaxy.
@@
∗ approximate NLTE corrections are calculated as an average for aluminum and independently for sodium and potassium (see text for details)
The [/Fe] ratio is an excellent indicator of a stream’s progenitor due to galactic chemical evolution. elements are primarily produced by core-collapse supernovae, while iron is produced by both core-collapse and Type Ia supernovae. The [/Fe] ratio thus starts high early then declines to lower values later as a galaxy forms. Efficiently star-forming galaxies with lower mass-loading factors like the Milky Way maintain a high [/Fe] until relatively high metallicities, while inefficiently star-forming dwarf galaxies with higher mass-loading factors have an downturn at lower metallicities (e.g., Tinsley, 1979; Matteucci & Brocato, 1990; McWilliam, 1997; Tolstoy et al., 2009; Sheffield et al., 2012). This downturn, often called an “knee”, tends to occur at lower metallicities in lower-mass galaxies (e.g., Tolstoy et al., 2009; Kirby et al., 2011, 2020; Frebel & Bromm, 2012; Helmi et al., 2018; Hill et al., 2019; Theler et al., 2020). Globular clusters can be understood in this context as inheriting the chemical patterns of their host galaxies at their birth metallicity, as they are the most massive star clusters to form in a galaxy (e.g., Forbes et al., 2018; Choksi et al., 2018; Reina-Campos et al., 2019). Thus, globular clusters born in the Milky Way will tend to be high-, while those born in dwarf galaxies will be low- (e.g., Mucciarelli et al., 2017; Usman et al., 2024).
Figure 8 shows that Leiptr’s stars are overall low-. The NLTE correction is negligible for magnesium (Mashonkina, 2013; Bergemann et al., 2017), calcium (Mashonkina et al., 2007; Spite et al., 2012), and titanium (Bergemann, 2011; Sitnova et al., 2016). While Milky Way halo stars and stars in Milky Way globular clusters M92 and NGC 2419 have [/Fe] (NGC 2419 has some low-magnesium stars, but this is from magnesium depletion due to multiple populations), the three lowest-metallicity Leiptr stars’ magnesium, calcium, and titanium abundances are within of [/Fe] = 0. Thus, the Milky Way halo and its globular clusters’ high- abundances are overall inconsistent with Leiptr. However, several dwarf galaxies have low [/Fe] ratios similar to Leiptr. At , lower-mass classical dwarf galaxies () like Draco, Ursa Minor, and Sextans, as well as the dwarf galaxy stream Elqui, have low [/Fe] abundances similar to Leiptr. The lowest-mass ultra-faint dwarf galaxies like Carina II, Reticulum II, and Horologium I () also mostly have low [/Fe] abundances at . In contrast, higher-mass dwarf galaxies like Sculptor, Fornax, and GSE () still have relatively high [/Fe] ratios at , not seeing a significant decline until higher metallicity, . Thus, Leiptr’s element abundances are most similar to lower-mass dwarf galaxies.
@@ 4.4 Odd-Z elements (Na, Al, K, Sc)
∗ approximate NLTE corrections are calculated independently for manganese (see text for details)
Figure 9 shows the abundances for the odd-Z elements. In blue points, we show the effect of applying corrections to sodium, aluminum, and potassium for NLTE effects on the lines where the radiation field is partially decoupled from local conditions. Using the calculations for neutral sodium from Lind et al. (2011), we compute the average of the corrections from each of the used lines for each star. All five stars include the 5895 Å sodium line and some of the five also use the 5682, 5688, and 5889 Å lines. The typical size of the correction is dex. Aluminum is determined using the 3961 Å resonance line, and we adopt a constant correction for all stars of adding 0.6 to [Al/Fe], following Roederer & Gnedin (2019). For potassium, we use the two resonance lines at 7664 Å and 7669 Å. The two lines always agree excellently in abundance, and we note the 7664 Å line is not affected by tellurics given Leiptr’s range of radial velocities (Figure 5). We identify stars in Table A.1 of Reggiani et al. (2019) that have the most comparable stellar parameters as the Leiptr stars and use the corresponding potassium NLTE corrections. We use the corrections from stars HD 186478, CS 29513-014, BD +19 1185A, HD 45282, and HD 126238 with a range from to dex. The NLTE correction is negligible for scandium (Zhang et al., 2014). The LTE and NLTE abundances are both reported in Tables 1-4.2.
In Figure 9, compared to halo stars (whose abundances are mostly computed in LTE and should be compared to the light-blue points), Leiptr lies on the lower end of the abundance scatter for sodium, potassium, and scandium. Abundances of aluminum in the Milky Way globular clusters, M92 and NGC 2419, are also substantially higher than in all four Leiptr stars. The globular clusters should be compared to the NLTE aluminum abundances (blue points): the M92 stars use the same aluminum 3961 Å line and were corrected using Nordlander & Lind (2017) while the NGC 2419 stars use the subordinate doublet at Å that are minimally affected by NLTE. These aluminum abundances further disprove that Leiptr may be an accreted globular cluster with multiple populations, which should result in aluminum enhancements as shown by the globular clusters in Figure 9 (see Bastian & Lardo, 2018; Gratton et al., 2019, for details).
@@ 4.5 Iron-peak elements (Cr, Mn, Co, Ni, Zn)
Figure 10 shows some of the iron-peak abundances in Leiptr. Manganese deserves special discussion due to NLTE effects. We reject the resonance lines at 4030 Å and only use weaker lines around 4800 Å. This was motivated due to expecting relatively small NLTE corrections (Bergemann & Gehren, 2008), but more recent calculations show that the corrections for the weaker lines can also be dex (Bergemann et al., 2019), so we apply an individual NLTE correction from Bergemann et al. (2019) to each star from the average of the corrections for each line, which appear as blue points in Figure 10. We note that the comparison stars in Figure 10 do not have NLTE corrections applied, so these should be compared to the LTE abundances. The NLTE correction is negligible for nickel (Eitner et al., 2023) and zinc (Takeda et al., 2005; Ezzeddine et al., 2019).